# Positive Effects of Financing

### v3

@@ 3 Discussion
We analyze the DICE integrated assessment model and derive metrics to assess intergenerational equity of the optimal emission pathways calibrated by the model. We extend the DICE integrated assessment model and compare several extensions for their effect on intergenerational equity. Our analyses apply to integrated assessment models in general and could also be extended to benefit-cost analysis of climate mitigation. We provide a complete open-source implementation of the model illustrating all our extensions, utilizing Monte-Carlo implementation of standard interest rate models for simulation and (stochastic) algorithmic differentiation to efficiently calculate sensitivities of model quantities against each other.
Our first observation is that the DICE model can significantly increase the policy cost measured in terms of cost-per-GDP. This is because the optimization does not consider any (cost-related) intergenerational equity but balances the marginal aggregated cost changes induced by a policy change. Absolute cost and its distribution over time are irrelevant to the objective function. This effect is driven by the aggregation of discounted welfare into intergenerational welfare.
Considering uncertainty in the rate of time-preference and thus overall interest rates, this effect is amplified when the stochastic adaptation of the abatement strategy generates a significant risk of increased cost. This indicates that a repeated re-calculation of the DICE model to current interest rate levels will amplify intergenerational inequality.
We consider natural model extensions by allowing the funding of abatement cost and non-linear discounting of damage cost. Funding of abatement cost may be considered as modeling the backing of abatement costs by government loans. We find that this may have a positive effect on intergenerational equity in terms of the cost distribution as it improves the alignment of abatement and damage cost. The non-linear discounting of damage cost may be considered as modeling financing risk of damage costs, where very large cost have higher financing cost. The non-linear damage cost can be used as a constrain that allows to effectively limit the generational cost to a specific percentage of the GDP.
@@
Modifying the DICE model in a modular way, we address the issue that the classical model does not consider uncertainty in interest rates and its effects on equity between generations. The presented extensions accounting for the funding cost of abatement and increased financing cost of large damages reduce intergenerational inequality. This is amplified when considering stochastic interest rate risk due to convexity induced by the DICE objective function. Constraining the total costs of mitigation and damages relative to GDP could be used as a simple objective to ensure intergenerational equity of climate mitigation.
@@ 4 Methods
The starting point of our study is the 2016 implementation of the classical DICE model [6, 7], since the DICE-2023 description [38] was only published parallel to this research.
Our implementation can deal with an arbitrary time-discretization $\{t_{i}\}$ and the DICE model is re-defined as being an Euler-discretisation of a time-continuous model. For the equidistant discretisation with $\Delta t_{i}=5\leavevmode\nobreak\ \text{years}$ the classical DICE-2016 model is recovered. Our numerical experiments are based on a time-discretization with $\Delta t_{i}=1\leavevmode\nobreak\ \text{year}$. The differences to the classical DICE-2016 are negligible for our work.
@@
Except where otherwise stated, we use the same model components (e.g. abatement cost and damage cost model) as in DICE-2016. A detailed description of our implementation of the DICE model and its extensions can be found in [40]. We use the same notation, except that we denote the cost functions by $C$ [^1] In [7] the letter $C$ was used for consumption, which we do not reference here.
We will first recapitulate the role of the time-preference rate $\rho$ in the DICE model in Section 4.1 as a background to our extension with stochastic time-preference rates.
Second, we recall the optimization procedure in Section 4.2, where we introduce a parameter-reduced form of the abatement policy to use for our analysis.
Third, we present metrics to analyze the temporal distribution of cost and the temporal distribution of cost-sensitivity to policy changes in the classical DICE model in Section 4.3.
Following upon this concepts, we present our extensions to the model, which change the distribution of cost in Section 4.4. We perform our analysis for the classical (deterministic) DICE model, as well as a model extended with stochastic interest rates and a stochastic abatement policy.
Finally, we introduce the concept of convexity from mathematical finance in section 4.5.
4.1 Time Preference Rate

### v4

@@ 4 Discussion
We analyzed the DICE integrated assessment model and derived metrics to assess the temporal cost structure and thus intergenerational equity of the optimal emission pathway. We extend the DICE integrated assessment model and compare several extensions for their effects on intergenerational equity. Our methods and extensions apply to integrated assessment models in general and can also be extended to the benefit-cost analysis of climate mitigation.
Our first observation is that the DICE model optimization results in an unequal temporal distribution of policy costs measured in terms of cost-per-GDP. This is because the optimization does not consider any (cost-related) intergenerational equity but balances the marginal aggregated cost changes induced by a policy change, while absolute cost and its distribution over time are irrelevant to the objective function.
Considering uncertainty in the rate of time-preference and thus overall interest rates, this effect is amplified when the stochastic adaptation of the abatement strategy generates a significant risk of increased cost. This indicates that repeated re-calculation of the DICE model to current interest rate levels amplifies intergenerational inequality.
We consider natural model extensions by allowing the funding of abatement cost and non-linear discounting of damage cost. Funding of abatement cost may be considered as modeling the backing of abatement costs by government loans. We find that this may have a positive effect on intergenerational equity in terms of the cost distribution as it improves the alignment of abatement and damage cost. The non-linear discounting of damage cost may be considered as modeling financing risk of damage costs, where very large cost have higher financing cost. The non-linear damage cost can be used as a constraint that allows to effectively limit the generational cost to a specific percentage of the GDP.
@@
Modifying the DICE model in a modular way, we address the issue that the classical model does not consider uncertainty in interest rates and its effects on equity between generations. The extensions presented accounting for the funding cost of abatement and increased financing cost of large damages reduce intergenerational inequality. This is amplified when considering stochastic interest rate risk due to convexity induced by the DICE objective function. Constraining the total costs of mitigation and damages relative to GDP could be used as a simple objective to ensure intergenerational equity of climate mitigation. Our results can also be interpreted as supporting the design of financing mechanisms such as a climate transformation fund that channels dedicated, long-term resources into mitigation and adaptation, see also [45], which builds on the modeling framework developed in the present work and verifies that a climate transformation fund can be funded with less than 3 % of the GDP. Implementing these mechanisms in practice would require collaboration between policymakers, financial institutions, and affected communities to ensure that they are economically sustainable, politically viable, and perceived as fair between generations.
@@ 5 Methods
The starting point of our modeling exercise is DICE-2016 [6, 7].
@@
Except where otherwise stated, we use the same model components as in DICE-2016. A detailed description of our implementation of the DICE model and its extensions can be found in [46]. We use the same notation, except that we denote the cost functions by CC. [^1] In [7] the letter $C$ was used for consumption, which we do not refer to here.
We redefine the objective function to be suitable for stochastic model extensions; see Section 5.1. We then analyze the temporal distribution of cost and the cost sensitivity to policy changes in the classical DICE model in Section 5.2 as an indicator of intergenerational equity. Section 5.3 introduces model extensions, which allow to alter the distribution of cost towards intergenerational equitable climate mitigation. Most importantly, we allow for a stochastic time preference $\rho$ modeled by a classical interest rate model, as described in section˜5.3.1.
5.1 Model Objective
The objective of the DICE model is to determine the abatement rate $t\mapsto\mu(t)$ and the savings rate $t\mapsto s(t)$ that maximize a given objective function.
The objective function of the classical DICE model is the integrated discounted welfare
$\int_{0}^{T}V(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t\text{.}$ (1)
Here, $V$ denotes the utility and $N$ denotes the numéraire. DICE-2016 is recovered for $N(t)=\exp(\rho\cdot t)$. We use the more general notion of a numéraire, as we will consider a more general model later. For the stochastic model, the objective function will be an expectation of (1) and a quantile for uncertainty analysis.
5.1.1 Abatement Policy - Classical DICE Model and Reduced Model
In the time-discrete classical DICE model, optimization determines the optimal parameters $i\mapsto\left(\mu(t_{i}),s(t_{i})\right)$.
For comparison with a parametric stochastic abatement model, which we introduce later as part of our model extensions, we define a parametric abatement model
where the parameter $T^{\mu=1}$ represents the time required to reach 100% abatement. This model has only a single parameter, however it turns out that the optimal abatement policy in the full parameter model is comparably close to a functional form of the type (2).
5.2 Metrics Describing Intergenerational Burden
To characterize the resulting optimal mitigation paths with respect to Intergenerational equity, we consider the total cost.
5.2.1 Cost
We analyze the temporal distribution of cost $t\mapsto C(t)$, which is the sum of the abatement cost $C_{\mathrm{A}}$ and the damage cost $C_{\mathrm{D}}$
$C(t)\ :=\ C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t)\text{,}$
and investigate the interplay of abatement and damage cost.
5.2.2 Cost per GDP
We measure the burden borne by a generation by considering the total cost per GDP, as it is common in other contexts, for example, to assess national debt [35]. Here, it can be considered as mimicking an effort-sharing scheme, that a more wealthy or productive generation can carry a greater burden,
To allow for a netting within a generation, one may consider the (discounted) average value over a generation lifetime. For a given time $s$ let $T^{\mathrm{L}}(s)$ denote the expected lifetime of a generation. We then consider the lifetime-average cost and GDP,
In our numerical experiments, we measure the instantaneous cost-per-GDP $\frac{C(s)}{GDP(s)}$ as well as the lifetime-average cost-per-GDP $\frac{\overline{C}(s)}{\overline{GDP}(s)}$.
5.2.3 Sensitivity of Damage and Abatement Cost to Policy Changes
The following shows why the unconstrained optimal emission path may result in an unequal distribution of the total cost and the total cost per GDP.
The optimal abatement policy $\mu$, representing the equilibrium state of the model, fulfills (differentiating the objective function (1) with respect to the abatement policy $\mu$)
[…]
As the objective function depends on the cost, two transformations take place. First, the value $V$ is defined as utility. Because the utility function is concave, the weight $\frac{\partial V(t)}{\partial C(t)}$ decays. Second, the values are discounted. With a constant positive time-preference rate, the discount factor constitutes an exponentially decaying weight $\frac{N(0)}{N(t)}$ applied to the cost. The product of both weights represents the cost-to-value-weight.
From eq.˜3, we see that the model optimization may result in some intergenerational inequality: Let $s$ be a fixed time at which a change in the abatement policy $\mu(s)$ is considered. Because $\mathrm{d}C_{\mathrm{D}}/\mathrm{d}\mu(s)<0$ and $\mathrm{d}C_{\mathrm{A}}/\mathrm{d}\mu(s)>0$ we see from eq.˜3 that the model equates the decrement in the cost of damage to the increment in the cost of abatement, both weighted by the value-to-cost-sensitivity $\mathrm{d}V(t)/\mathrm{d}C(t)\cdot N(0)/N(t)$. We have
$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ -\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\text{,}$ (4)
where $t\mapsto\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}$ and $t\mapsto\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}$ exhibit different temporal distributions.
The sensitivity $\frac{\mathrm{d}}{\mathrm{d}\mu}$ is a functional derivative or gradient. Given the one-parameter abatement model (2) the equilibrium state is given by
$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ -\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\text{,}$ (5)
and we can plot the (weighted) damage cost sensitivity $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ and the (weighted) abatement cost sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$, see fig.˜2.
5.2.4 Damage Cost per Abatement Cost upon Abatement Policy Change
A change in the abatement policy at time $s$ affects the abatement cost $C_{\mathrm{A}}(s)$ only at time $s$, whereas it affects damage cost $C_{\mathrm{D}}(t)$ for all $t\geq s$. This motivates to consider the damage cost per abatement cost induced by a change in the abatement policy $\mu(s)$, i.e.,
$\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\big{/}\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}\text{.}$
As the model optimization weights cost by the cost-to-value weight, we consider the corresponding weighted damage cost per abatement cost sensitivity.
Definition (Damage Cost per Abatement Cost upon Abatement Policy Change)
We define [^2] The minus in (6) accounts for the fact that abatement cost sensitivity and damage cost sensitivity have opposite signs.
$\gamma_{\mathrm{D}}(s;t)\ :=\ -\frac{\frac{\partial V(t)}{\partial C(t)}\cdot\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\cdot\frac{N(0)}{N(t)}}{\frac{\partial V(s)}{\partial C(s)}\cdot\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}\cdot\frac{N(0)}{N(s)}}\text{.}$ (6)
Lemma
At the optimal abatement policy $\mu$, for fixed $s$ the integral over $t\mapsto\gamma_{\mathrm{D}}(s;t)$ equals $1$, i.e., $\gamma_{\mathrm{D}}(s)$ can be interpreted as a density, and we may interpret
$t_{D}(s)\ :=\ \int_{0}^{\infty}t\cdot\gamma_{\mathrm{D}}(s;t)\ \mathrm{d}t\text{.}$
as the expected time of damage cost per avoided abatement cost at time $s$.
The proof is provided in the Appendix B.
Our numerical analysis will show a significant (expected) time lag for $t_{D}(s)$ (roughly 70 years in a DICE-2016). This means that marginal gains of one generation are equated to marginal losses of another generation. The optimization does not equalize the burden but rather the weighted marginal burden. In this sense, the optimization is indifferent to the temporal distribution of the absolute burden.
5.3 Model Extensions
We present modifications to integrated assessment models to investigate the effects of stochastic interest rates, funding abatement costs, and non-linear financing costs [47]. These may also be incorporated into more complex IAMs. We use the DICE model to illustrate potential effects.
5.3.1 Stochastic Interest Rates
In classical IAMs, the time preference rate $\rho$ is typically a time-dependent deterministic function, which is often a constant.
We model a stochastic interest rate (time-preference rate) $\rho$. The model provides the stochastic numéraire $N$ (a bank account accruing continuously at rate $\rho$) as well as the stochastic forward rate $FR$. Setting the volatility of the interest rate to zero recovers the deterministic model.
Our implementation allows for the use of a general discrete forward rate model (LIBOR Market Model). Our results were obtained using the classical Hull-White model [48, 49], that is
$\begin{split}\mathrm{d}\rho(t)&\ =\ (\theta(t)-a(t)\rho(t))\ \mathrm{d}t\ +\ \sigma(t)\ \mathrm{d}W(t)\text{,}\quad\rho(t_{0})\ :=\ r_{0}\text{,}\\ \mathrm{d}N(t)&\ =\ \rho(t)\ N(t)\ \mathrm{d}t\text{,}\end{split}$ (7)
We set $r_{0}$, $\theta(t)$ so that the certainty-equivalent rate $\bar{\rho}(t)$ equals the deterministic time preference rate $\rho$, such that outcomes are exactly identical to the deterministic model (9). Mean reversion and volatility were chosen to reflect plausible long-term interest-rate dynamics. The chosen mean-reversion speed of $a(t)=0.02$ implies a half-life of $\ln(2)/0.02\approx 35$ years, while the chosen volatility of $\sigma(t)=0.3\%$ produces a dispersion of future rates that remains within a realistic range—roughly $\pm 3$ percentage points around the long-run mean over a century.
While our result should be interpreted as illustrative rather than predictive, showing that stochastic interest rates generate adverse effects, the interest rate scenarios generated by our parameter choice are realistic, rather conservative: the 1st and 99th percentile of the generated interest rates at t = 100 years are -1 % and 5 %, respectively. For a corresponding visualization, see [46].
As we consider a model extension where $N$ and $V$ are stochastic processes, we redefine the objective function (1) as a random variable. For the optimization we consider the expected welfare
$\mathrm{E}\left(\int_{0}^{T}V(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t\right)\text{.}$ (8)
If interest rates are stochastic but the utility function $V(t)$ remains deterministic, adding stochastic interest rates does not change the interaction with the IAM because
$\begin{split}\mathrm{E}\left(\int_{0}^{T}V(t)\frac{N(0)}{N(t)}\mathrm{d}t\right)&\ =\ \int_{0}^{T}V(t)\ \mathrm{E}\left(\frac{N(0)}{N(t)}\right)\mathrm{d}t\\ &\ =\ \int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\bar{r}(s)\mathrm{d}s\right)\mathrm{d}t\text{,}\end{split}$ (9)
where $\bar{r}(t)\ =\ -\frac{\partial}{\partial t}\log\left(\mathrm{E}\left(\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\right)\right)$.
Owing to the lack of feedback that allows adjustment of the abatement policy and resulting damages, adding stochastic interest rates does not yet introduce changes in the model dynamics. The model produces the same emission pathway as for the deterministic model. However, this changes if one assumes that the abatement policy is adapted to the (stochastic) changes in the interest rate level.
5.3.2 Stochastic Abatement Policy
Assuming that a change in the interest rates triggers a re-calculation of the model optimal abatement policy, a stochastic interest rate implies a stochastic abatement policy. The abatement policy $t\mapsto\mu(t)$ becomes a stochastic process that adapts to changes in interest rates. This reflects the possibility that the abatement policy can be adjusted to the interest rate scenarios.
Determining an optimal stochastic abatement policy is a classical optimal exercise problem, as in the valuation of financial products with early exercise options in mathematical finance. Here, $\mu$ takes the role of the exercise strategy. As the time-$t$ optimal exercise is required to be a $\mathcal{F}_{t}$-measurable, it may be represented as a functional form of time-$t$ measurable random variables [49].
An approach to determine the optimal exercise strategy is to parameterize $\mu(t)$ as a functional form of $\mathcal{F}_{t}$-measurable random variables and apply a global optimization. As we utilize a one-factor Markovian model for $\rho$, we consider a functional form where $\mu$ depends on $t$ and the current interest rate level $\rho(t)$.
To obtain a robust optimization, we propose the parametric stochastic abatement model
The model can be seen as a first-order approximation of general policies. Its simplicity allows the interpretation of the parameters: $a_{0}$ is the abatement speed at the zero interest rate level and $a_{1}$ describes the change in the abatement speed per interest rate change,
The optimization of the model (10) results in a negative value for $a_{1}$, reflecting that a higher interest rate corresponds to a slower abatement speed.
5.3.3 Funding of Abatement Costs
The simple damage function of the DICE model provides damages as a reduction of global GDP. Damage may be reduced by abatement of emissions. Both abatement and damage are associated with costs. The time-$t_{i}$ costs are directly deduced from the time-$t_{i}$ GDP. The remainder is then available for consumption or investment. Investment adds to the capital, which determines the GDP of the next time $t_{i+1}$. However, abatement cost and damage cost resulting from the same abatement policy decision occur at significantly different times, as our analysis shows.
To resolve the time-lag between abatement and damage costs, we introduce the option of funding abatement cost. As the abatement of emissions is a planned process of societal relevance, it is reasonable to assume that abatement costs are covered by a loan for which interest corresponds to the current discount rate. Our model allows the application of a funding spread, i. e. higher interest rates for loans than for discounting, but this is not considered in this general introduction of the phenomenon. In future work, the concept of the climate beta could be used to introduce an interest rate spread for abatement investments comparable to the modifications in DICE-2023 [8].
We define $C_{\mu}(t)$ as the instantaneous abatement costs in time $t$, i.e., the quantity that was formerly denoted by $C_{\mathrm{A}}(t)$. We allow that $C_{\mu}(t)$ is funded for a period $\Delta T_{\mathrm{A}}$. Thus, the abatement costs of $C_{\mu}(t)$ are accrued with the forward rate $FR(t,t+\Delta T_{\mathrm{A}};t)$ observed in $t$, such that the realized abatement cost $C_{\mathrm{A}}$ at time $t+\Delta T_{\mathrm{A}}$ is
Instead of a single (zero-bond like) funding, we can also distribute the time $T_{i}$ cost $C_{\mu}(T_{i})$ to a collection of $m_{i}$ equal value loans with repayments in $T^{k}_{i}$, $k=i+1,\ldots i+m_{i}$. In this case the $T_{j}$ cost become
$C_{\mathrm{A}}(T_{j})\ :=\sum_{i\ :\ \exists\ T^{k}_{i}=T_{j}}\ \frac{1}{m_{i}}\ C_{\mu}(T_{i})\ \left(1+FR(T_{i},T_{j};T_{i})\ (T_{j}-T_{i})\right)\text{.}$ (12)
From a valuation (discounting) perspective, both funding approaches have no effect, as the accrual is compensated by discounting. However, the change in timing might introduce an effect in the DICE model due to the way abatement and damage costs are associated and temporal differences in each generation’s utility or capacity.
For a fixed emission pathway, the modification leaves present-value abatement costs unchanged (accrual cancels with discounting). The equity effect comes from changes in timing, not value creation.
In our numerical experiments, we distributed the funding according to (12) over 60 years ($m_{i}=60$).
5.3.4 Non-Linear Financing Costs
The present value of an unsecured financial cash-flow is defined by a discount factor times the cash-flow. If the future cash-flow is subject to default, the discount factor is lower, reflecting an additional value reduction due to the risk of (partial) default. As default is not an option for future damage costs, it appears as if the risk-free discount factor should apply. However, since no hedging strategy exists, a risk free funding is not possible. Thus, additional costs may be incurred to secure the unsecured funding [47]. Since damage cost may become very large - much larger than funds provided by standard financial markets - it is reasonable to assume that these additional funding cost become (over-proportionally) large for larger damages. We model this by optionally adding non-linear financing cost using a non-linear funding model [47].
This model allows the discount factor to depend on the magnitude of the cash-flow. Thus the funding of larger cash-flows requires a premium to compensate for a larger default risk or other frictions. For our application, this means that larger damages get a larger weight. Consequently, the abatement policy optimization will prefer emissions paths that avoid very large damages, which is beneficial with respect to intergenerational equity.
We modify the damage cost function. Let $C^{\circ}_{\mathrm{D}}(t)$ denote the damage cost of the classical model, i.e. the quantity that was formerly denoted by $C_{\mathrm{D}}(t)$. We then redefine the effective damage costs as
where $DC(C^{\circ}_{\mathrm{D}}(t);t)$ is the default compensation factor modeling over-proportional cost of funding large projects, [47]. The factor $DC$ is equal to 1 for modest damage values and larger for large damage values.
As $C^{\circ}_{\mathrm{D}}(t)$ represents a time-value, it is natural that the default compensation factor depends on a normalized value only, e.g., percentage of the GDP. In our numerical experiment we used $DC(C^{\circ}_{\mathrm{D}}(t);t)\ :=\ DC^{*}(\frac{C^{\circ}_{\mathrm{D}}(t)}{GDP(t)})$ with $DC^{*}(x)=10\%$ for $x\geq 3\%$ and $DC^{*}(x)=1$ otherwise, basically limiting costs to 3 % of GDP.
5.4 Implementation Details
The numerical experiments were conducted with an annual time-discretization $\Delta t_{i}=1$, using a Monte-Carlo simulation with 10000 sample paths. Sensitivities were obtained through adjoint automatic differentiation. Optimization was performed with an ADAM optimizer.
Supplementary Material
Supplementary material is available online.
Acknowledgements
Author Contributions
Competing Interests
The authors declare that they have no competing interests.
[…]
Supplementary information for
Intergenerational Equitable Climate Change Mitigation:
Negative Effects of Stochastic Interest Rates and;
Positive Effects of Financing
Christian P. Fries and Lennart Quante
Contents
[…]
[supplement]l1
Appendix A Supplementary Figures
We provide some additional figures. LABEL:fig:cost_abatement_damage_disounted shows the temporal distribution of the cost, which should be compared to the temporal distribution of the cost sensitivity in fig.˜2.
LABEL:fig:pathways_deterministic_with_extension and LABEL:fig:pathways_stochastic_with_extension show the associated model quantities (emission, abatement, damage, and total cost per GDP) for the deterministic and stochastic model, with and without our model extensions.
LABEL:fig:opt_abatement_time_by_interest_level and LABEL:fig:opt_abatement_time_by_interest_volatility show the sensitivity of the abatement policy on the interest rate level and interest rate volatility.
[…]
figureList of Supplementary Figures
Supplementary Figure 1: The abatement cost $C_{\mathrm{A}}$ (green) and damage cost $C_{\mathrm{D}}$ (red), discounted by $\frac{N(0)}{N(t)}$, in the original DICE model. The optimization is not balancing the temporal distribution of the cost - which is ultimately resulting in an intergenerational inequality.
supfigure]fig:cost_abatement_damage_disounted
Supplementary Figure 2: Deterministic model: Financing extensions lead to faster abatement compared to standard model; a emissions, b abatement, c damages, and d total GDP relative cost.
supfigure]fig:pathways_deterministic_with_extension
Supplementary Figure 3: Stochastic model: Financing extensions lead to faster abatement compared to standard model; a emissions, b abatement, c damages, and d total GDP relative cost. Shaded area shows the 10th to 90th percentiles (VaR).
supfigure]fig:pathways_stochastic_with_extension
Supplementary Figure 4: Stochastic DICE Model with Different Interest Rate Volatilities ($\sigma=0.2\%$ (green), $\sigma=0.3\%$ (blue), $\sigma=0.4\%$ (red)). Allowing adaptation of the abatement policy to stochastic interest rates increases intergenerational inequality and exhibits a significant risk of even higher intergenerational inequality (a: emissions, b: abatement, c: damages, and d: total GDP relative cost). Shaded area shows the 10th to 90th percentiles.
supfigure]fig:pathways_stochastic_variation
supfigure]fig:opt_abatement_time_by_interest_level
supfigure]fig:opt_abatement_time_by_interest_volatility
Appendix B Damage Cost per Abatement Cost upon Abatement Policy Change
We show that the damage cost per abatement cost upon abatement policy change is a density if $\mu$ is in its equilibrium state.
If $\mu$ denotes the optimal abatement policy we have
[…]
For the time discrete model where $C_{\mathrm{A}}(s)$ depends on $\mu(s)$ only, we find at $s=t_{i}$ that
$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ =\ \frac{\mathrm{d}V(t_{i})}{\mathrm{d}C(t_{i})}\frac{\mathrm{d}C_{\mathrm{A}}(t_{i})}{\mathrm{d}\mu(t_{i})}\ \frac{N(0)}{N(t_{i})}\ \Delta t_{i}\text{.}$
Hence, for $\Delta t_{i}=1$ (using $s=t_{i}$)
$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ =\ -\frac{\mathrm{d}V(s)}{\mathrm{d}C(s)}\frac{\mathrm{d}C_{\mathrm{A}}(s)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(s)}\text{.}$
and thus
$\int_{0}^{T}-\frac{\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}}{\frac{\mathrm{d}V(s)}{\mathrm{d}C(s)}\frac{\mathrm{d}C_{\mathrm{A}}(s)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(s)}}\ \mathrm{d}t\ =\ 1\text{.}$
Appendix C Non-Linear Financing Costs
We modify the damage cost function as
where $C^{\circ}_{\mathrm{D}}(t)$ denote the damage cost of the classical model and $DC(C^{\circ}_{\mathrm{D}}(t);t)$ the default compensation factor modeling the over-proportional cost of funding large projects, [47].
We allow for two different normalizations: with the numéraire $N(t)$ or with the GDP $Y(t)$. Given a function $DC^{*}$, the model for the default compensator is
$DC\left(C^{\circ}_{\mathrm{D}}(t);t\right)\ =\ DC^{N}\left(C^{\circ}_{\mathrm{D}}(t);t\right)\ :=\ DC^{*}\left(\frac{C^{\circ}_{\mathrm{D}}(t)}{N(t)}\right)\text{,}$ (13)
or, alternatively,
$DC\left(C^{\circ}_{\mathrm{D}}(t);t\right)\ =\ DC^{Y}\left(C^{\circ}_{\mathrm{D}}(t);t\right)\ :=\ DC^{*}\left(\frac{C^{\circ}_{\mathrm{D}}(t)}{Y(t)}\right)\text{.}$ (14)
The factor $DC^{*}(x)$ depends on the size of $x$. For small $x$ we have $DC^{*}(x)=1$, but for large $x$ we may have factors >1>1. Obviously, this will penalize large spikes in the costs.
Appendix D Model Implementation
We provide a complete open-source implementation [51] of the model illustrating all our extensions, utilizing the Monte-Carlo implementation of standard interest rate models for simulation and (stochastic) algorithmic differentiation to efficiently calculate the sensitivities of model quantities against each other.
Our implementation can deal with an arbitrary time-discretization $\{t_{i}\}$ and the DICE model is redefined as the Euler-discretization of a time-continuous model. For the equidistant discretisation with $\Delta t_{i}=5\leavevmode\nobreak\ \text{years}$ the classical DICE-2016 model is recovered. Our numerical experiments are based on a time-discretization with $\Delta t_{i}=1\leavevmode\nobreak\ \text{year}$, however, the differences to the $\Delta t_{i}=5\leavevmode\nobreak\ \text{years}$ discretization of the classical DICE-2016 are negligible in our study. With annual time-discretization the piecewise constant abatement policy constitutes $499$ free parameters for a $T=500\leavevmode\nobreak\ \text{years}$ model.
In the numerical experiments an annual time-discretization $\Delta t_{i}=1$ was
Appendix E Associated Concepts
E.1 Time Preference Rates and the Ramsey Equation